What is the value of y in the solution to the following system of equations?
5x – 3y = –11
2x – 6y = –14

Found 2 solutions by Alan3354, Drum3freak:
Answer by Alan3354(69443)

5x – 3y = –11
2x – 6y = –14
------------ Multiply eqn 1 by 2
10x –6y = –22
2x – 6y = –14
-------------- Subtract
8x = -8
Can you finish it?

Answer by Drum3freak(1)

Solve one of the two equations for x in terms of y. The first equation 5x-3y=-11 would be -3y=-11-5x --> -y=-(11/3)-(5/3)x --> y=(11/3)+(5/3)x. now for the other equation plug this equation in for y and solve for x, from there you will have only one variable and you can solve for y.